a NEW CAR SALESMEN KNOWS THAT HE SELLS CARS TO ONE CUSTOMER OUT OF TEN WHO ENTERS THE SHOW ROOM.WHAT IS THE PROBABILITY THAT HE WILL SELL A CAR TO EXACTLY 2 OF THE NEXT 3 CUSTOMERS.IS IT:
1)0.027
2)0.973
3)0.000
4)0.090
5)0.901

Answer by stanbon(75887)

a NEW CAR SALESMEN KNOWS THAT HE SELLS CARS TO ONE CUSTOMER OUT OF TEN WHO ENTERS THE SHOW ROOM.WHAT IS THE PROBABILITY THAT HE WILL SELL A CAR TO EXACTLY 2 OF THE NEXT 3 CUSTOMERS.IS IT:
-----
Binomial Problem with n=3 ; p = 0.1 ; x = 2
P(x=2) = 3C2(0.1)^2*(0.9) = 3*0.01*0.9 = 0.027
=============
Cheers,
Stan H.
=============

1)0.027
2)0.973
3)0.000
4)0.090
5)0.901